OpenStudy (anonymous):

The population of a region is 100 000. There are 250 albinos in the region. The estimated number of normal person carrying albino gene is: a. 5900 b. 9500 c. 95000 d.59000 e. 159000

OpenStudy (anonymous):

To answer this you need background info on the type of albinoism. Is it a single gene? Is it dom or res? Have to know to answer.

OpenStudy (anonymous):

Assuming the type of albinoism is ressesive and monogenic, which it almost always is. Assuming the population is at hardy-weinberg equilibrium, which you need to answer the question. q^2+2pq+p^2=1 (the hardy-weinberg equilibrium) Where q^2 is the number of people with albinism, 2pq is the number of people heterozygotic for the trait and p^2 is the number of homozygotes for the healthy gene. Where q^2=250/100000=0.0025 q=0.05 p=1-0.05=0.95 2pq=2*0.05*0.95=0.095 0.0475*100 000=9500

OpenStudy (anonymous):

So the answer is b

OpenStudy (anonymous):

@niels5x9 thankss...you too thankss @mrdoldum